Which of the following statements are true or false? Justify your answer
with appropriate explanations
a) An increase in the price of ham decreases the consumption of cheese.
An increase in the price of ham increases the consumption of salami.
Hence ham and cheese are substitutes and ham and salami are
complements.
b) If price elasticity of demand (PED) is -1.7, it is feasible that a company
cuts the price by 5% in order to generate enough total revenue
1
Expert's answer
2016-03-18T15:48:04-0400
a) according to the definition a complement or a complementary good is a good which demand is increased when the price of another good is decreased. A substitute is a good which demand is increased when the price of another good is increased.
Ham and cheese are complements and ham and salami are substitutes by definition.
Therefore, the statement a) is false.
b) price elasticity of demand is defined by the following statement: Ed = (dQ/Q) / (dP/P)
Which means that if the company cuts the price by 5%, demand will rise by 8,5%. So before the price cut the total revenue was PQ, and becomes (1-5%)P * (1+8.5%)Q = 0.95P*1.085Q = 1.03075PQ after. So by cutting the price by 5% the company increases the total revenue by 3.075% Generally speaking, it is not clear from the condition of the task what is "enough total revenue" exactly - i.e. what target levels the company has, but since the price cut leads to positive impact on the revenue, we may say that the statement b) is true, that is, it is feasible that the company cuts the price by 5%